Parth (parthkohli):

Suppose that: 1 + 1 = 7 2 + 2 = 89 3 + 3 = 228 4 + 4 = ?

Parth (parthkohli):

Let's see how good you are at arithmetic. >:-)

Parth (parthkohli):

What kind of module?

OpenStudy (shubhamsrg):

404 ? (going by a forced logic !) -_-

OpenStudy (opcode):

520 is Opcode guess?

Parth (parthkohli):

@shubhamsrg @Opcode no

OpenStudy (anonymous):

8!!!!

Parth (parthkohli):

That's right! ^

Parth (parthkohli):

Does no one know 4 + 4 = 8?!

OpenStudy (anonymous):

for realz

OpenStudy (anonymous):

What about 1 + 1, 2 + 2, and 3+3? Or is it just a trick?

Parth (parthkohli):

@RD⁴² ?

Parth (parthkohli):

What about them?

OpenStudy (anonymous):

1 + 1 = 7 2 + 2 = 89 3 + 3 = 228 Why?

OpenStudy (anonymous):

I don't get the joke, honestly.

Parth (parthkohli):

I have asked you to just assume that. I didn't ask you to get to a conclusion that 1 + 1 is related to 4 + 4.

OpenStudy (anonymous):

It makes no sense, but ok.

OpenStudy (anonymous):

>: )

OpenStudy (shubhamsrg):

Here was my forced logic : 7= (4)(5)/2 -3 89 = (13)(14)/2 -2 228 = (21)(22)/2 -3 So for odd terms like 1+1, or 3+3 , we subtract 3 from n(n+1)/2 for even terms, we subtract 2 now for n, it is 4,13,21 --> So from the differences, I say the next term is 28 hence 4+4 = (28)(29)/2 -2 =404 :3

Parth (parthkohli):

Sir, you have my respect.

OpenStudy (opcode):

Too lazy to show my work :-). But I gave him a medal because he did ^_^.

OpenStudy (shubhamsrg):

LAWL !!

OpenStudy (anonymous):

I know this is late, but I object. It's 424. Let \(\large x + x = \dfrac{1}{2}(57x^2 - 7x - 36)\) When x = 1, 1 + 1 = 7 When x = 2, 2 + 2 = 89 When x = 3, 3 + 3 = 228 Lastly, when x = 4, 4 + 4 = \(\boxed{424}\) http://www.wolframalpha.com/input/?i=table [%2857+x^2+-+7x+-+36%29%2F2%2C+{x%2C1%2C4}] Looks like it is safe to say that this question is ambiguous. Sir, you are not supposed to post an ambiguous question, @ParthKohli

Parth (parthkohli):

how is this ambiguous? do you not know 4 + 4 = 8?

OpenStudy (anonymous):

Yet, you said 1 + 1 = 7

Parth (parthkohli):

I asked you to suppose 1 + 1 = 7. I didn't ask you to assume that 4 + 4 will be based on 1 +1, 2 + 2 and 3 + 3.

OpenStudy (anonymous):

It does since 2 + 2 = \(\large 8\)9 and 3 + 3 = 22\(\large 8\) Beside shubhamsrg has other answers. There are several reasonable answers for this, hence ambiguous.

OpenStudy (anonymous):

Also, 'suppose' means 'assume,' according to dictionary. http://dictionary.reference.com/browse/suppose

OpenStudy (anonymous):

@micahwood50 Thank you! @ParthKohli He has good points...

OpenStudy (opcode):

It's a trick question :3 We fell for it :P

OpenStudy (anonymous):

Oh and don't you know that 1+1 ACTUALLY equal 2?

OpenStudy (anonymous):

If 1+1 =7, then it affects numeric system, which can also affect the value of 4.

OpenStudy (anonymous):

And we know that 2+2 =89 so if 4+4 = 8 then 2 is bigger than 4.

OpenStudy (anonymous):

You have not given whether numeric system affects 4 or not, hence ambiguous.

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

If I haven't given so, you shouldn't assume that 4 + 4 is based on the prior sequence.

OpenStudy (anonymous):

so you admit this question is ambiguous